Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 12

1. Section 5.6, problem 115:

Use the hint to find that I + I = ∫_{0}^{a} 1dx.

2. Chap 5, Chap 5, Additional and Advanced Exercises: problems 5 There are 2 solutions. x = 2 gives f(t) = \frac{1}{4}, x = −2 gives f(t) = \frac{−1}{4}. One can check that both give currect anxwer after substituting in.